OpenStudy (anonymous):

Using elimination to solve the system 2x + y = 8 and x + y = 4, which of the following could be the first step? A. 2x + y = 8--------> 2x + y = 8 (x + y = 4)(-1)---> -x - y = -4 B. 2x + (4 - x) = 8 C. 2x + y = x + y D. 3x + 2y = 12 E. x = 4

OpenStudy (anonymous):

D. Lol, this is so easy.

OpenStudy (anonymous):

really i am really bad with these

OpenStudy (anonymous):

All you do is take 2x and x add those together, then add the y's, then add the = stuff. That's how you start solving it.

OpenStudy (anonymous):

Solve the system. What is the value of x? y = 3x and 4x + 2y = 30 A. 3 B. -3 C. 9 D. -9 and i don't get these either

hero (hero):

A should be it

OpenStudy (anonymous):

2x+y=8 x+y=4 subtract. x=4 i'd say b